# Bribe & Fork: Cheap PCN Bribing Attacks via Forking Threat

### v1

@@ 2.2 Bribe & Fork Attack
Figure 1: Comparison of the scenarios: (a) Honest execution: once an old state appears on-chain (black rectangle), $\mathcal{P}_{1}$ gets an option to revoke this state with a transaction $tx_{1}$ (included in the block $txs_{1}$ which is published in the first round). (b) Attack in [10]: the bribing party publishes $tx_{2}$ and $tx_{b}$ included in a single block $txs_{2}$, with a large miner fee (reversely proportional to the fraction of the mining power $\lambda_{min}$ of the least significant miner). The miners skip mining $txs_{1}$ in the first round, and mine $txs_{2}$ in the last round. (c) Bribe & Fork: the bribing party publishes $txs_{2}$ with a fee sufficient to bribe only the strongest miner (with $f_{2}-f$ reversely proportional to $\lambda_{s}$). The strongest miner publishes the self-penalty transactions $tx_{p_{1}},tx_{p_{2}}$ that can be mined in transaction sets $txs_{p_{1}},txs_{p_{2}}$. In the first round, the miner $N_{s}$ locks $P\approx\lambda_{s}\cdot B$ to the deposit transaction $tx_{p_{1}}$, thus threatening other miners that they will be forked once $txs_{1}$ is mined before the deadline. After the deadline the transaction set $txs_{2}$ is published and the miner $N_{s}$ may collect back the deposit using $txs_{p_{2}}$.
@@ 3 Our Model
In this section, we introduce our model of PoW mining that takes into account the forks. First, we present the assumptions of our model and then introduce a set of new notions; a summary of our notation can be found in Table 1 (Appendix 0.B). Finally, we define the game that models the process of mining.
3.1 Assumptions
We introduce the assumptions that allow us to systematically study the behavior of the players who take part in PoW mining process.
Average transactions. In Section 0.D.1, based on empirical data, we show how block rewards fluctuate in the real-world. However, following [10], we assume that standard transactions have a constant (average) reward and that all blocks have a constant number of transactions. We refer to each standard transaction as an average transaction.
Abandon rules for forked states. In general, miners have the option to fork a chain or mine one of the existing chains from an arbitrary block of the chain. However, forking that starts at some old block is unrealistic; hence, we exclude this possibility from the game (following [24, 34, 29]). In other words, miners in our model can fork only the last transaction on one of the chains, and then either the original chain or the fork becomes stable whenever it reaches a length equal to the length of the original chain plus one. The forks can be modelled differently, e.g. assuming a longer fork length or using a finite automaton definition. We expect our results to hold in the alternative modelling as well, although the parameters of our solution could change.
PoW Mining. We focus on proof-of-work blockchains that employ the so-called Nakamoto consensus extending the longest chain, such as Bitcoin [27]. We assume that in such environments miners (players) tend to form mining pools to (a) bypass the task of verifying transactions – the pool’s manager dispatches a list of valid transactions for inclusion – and more importantly, (b) to guarantee a more stable income as individual mining carries substantial deviation. Empirical evidence supporting this assumption, drawn from Bitcoin, will be detailed in Section 0.D. For the rest of this work, each miner is assumed to either mine independently or stick to a selected mining pool for the whole course of the game, i.e., we will treat the mining pools as single players in the game.
Furthermore, in our analysis, we assume that typically there exists a miner that has a mining power of at least $20\%$ of the total hash power. This assists the execution of the feather-fork attack that requires a credible threat, i.e., a threat by a player with relatively strong mining power. Moreover, we assume that there exists at least one other player in the game that has a mining power slightly higher than $1\%$ and that no players with overwhelmingly small mining power (less than a $10^{-100}$ fraction of the total mining power) participate in the game. We back these assumptions with a detailed analysis of the distribution of miners’ computing power on real-life data in Section 0.D.2.
Preventing shallow forks. In Conjecture 1, we assume that players will have the option to fork some transaction only when they see an explicit opportunity of mining some transaction in the future, initially blocked by the currently forked transaction. In other words, we assume that players will fork what we call “average” transactions. That is, we forbid shallow forks in the model. This is because, in the real-world, excessive forking can trigger a denial of service attack on the blockchain. Thus, it is very unlikely that miners will follow the strategy of forking arbitrary transactions.
3.2 Preliminaries
Let us begin by recalling the conditionally-timelocked output definition from [10].
Definition 1 (Conditionally timelocked output - quoted from [10])
@@ 3.2 Preliminaries
For instance, let $\mathcal{S}=\{\ [(\mathsf{block}_{1},W_{1}),(\mathsf{block}_{2},W_{2})]\ \}$ be a current state with a single chain $S_{1}$. Then, after a successful fork of $S_{1}$ with $(B_{3},W_{3})$, one gets $\mathcal{S}=\{\ [(\mathsf{block}_{1},W_{1}),(\mathsf{block}_{2},W_{2})],[(\mathsf{block}_{1},W_{1}),(\mathsf{block}_{3},W_{3})]\ \}$.
Types of Transaction Sets. Each block mined in the game (denoted as $\mathsf{block}$) includes only one of the transaction sets listed below. Further, the block can be appended to the blockchain, whenever all of the following conditions are met:
@@
  • A (bribing) transaction set $txs_{2}$ that contains the transaction $tx_{2}$ that spends money of $txo$ under $cond_{2}$ and a bribing transaction $tx_{b}$. As long as $txo$ is not spent, this transaction set can be mined on a chain of length $\geq T$. The rest of the transactions in this transaction set are unrelated transactions $tx_{u}$.
@@
Rewards. We assume that a miner, after successfully mining a transaction set $txs_{i}$ on the main chain, gets a reward $\mathsf{reward}(txs_{i})$ equal to $B+f_{i}+P$, where $B$ is a constant block reward and $f_{i}$ is a sum of user fees input by users posting transactions in the transaction set $txs_{i}$. Whenever $txs_{i}$ contains a transaction which locks C coins from miner’s account, we set P to be equal to $-C$. Analogously, when a miner collects $C$ as one of the transactions from $txs_{i}$, we set the parameter $P$ to $C$. The reward for mining a block depends on the number of transactions within the block and their fees. The fee for a more complex transaction is typically higher as it occupies more space in a block. In this respect, we make the following assumptions that correspond to the current Bitcoin values (see Section 0.D.1 for more details):
@@
Mining Power Distribution. We assume the following mining power allocation among the players (see Section 0.D.2), denoted $\lambda=(\lambda_{1},\ldots,\lambda_{n})$:
@@
The Abandon Rule. Let us define the abandon rule $\mathsf{abandon}:\mathbb{S}\rightarrow\mathbb{S}$ that makes a chain the only chain in the game whenever it outruns the length of other chains in the game. That is, for any $S_{i}\in\mathcal{S}$: $\exists_{S_{j}\in\mathcal{S}}\ len(S_{j})\geq len(S_{i})+1$ then the abandon rule removes $S_{i}$ from the state $\mathcal{S}$.
No Shallow Forks Conjecture. We conjecture that players will not start shallow forks in the game. Shallow forks occur whenever miners fork a transaction that does not explicitly block an opportunity to mine some other transaction in the future.
Conjecture 1 (No shallow forks)
@@ 3.3 The Game
Finally, we describe a game that models the process of Proof-of-Work blockchain mining taking into account the option to conduct a forking of a block. The game proceeds in rounds, in each round the miners are able to choose whether they want to continue mining one of the chains or they want to fork one of the chains.
@@
When the final round $R$ of the game is over, it finishes in some state $\mathcal{S}$ and rewards are given to the players. By $\mathcal{S}^{*}$ let us denote the longest chain in the state $\mathcal{S}$. Whenever the state has multiple longest chains, $\mathcal{S}^{*}$ denotes the oldest of the longest chains of $\mathcal{S}$. After the final round $R$, in the state $\mathcal{S}$, the reward given to a player $N_{i}$ is: $\mathsf{reward}_{i}(\mathcal{S})=\sum_{(\mathsf{block},W)\in\mathcal{S}^{*}:W=i}\mathsf{reward}(\mathsf{block}).$
@@
4 Analysis
Here, we present a summary of our analysis that can be found in Appendix 0.C.
We prove that the Bribe & Fork attack is effective in a sequence of steps. We begin by showing that once the dispute period is over (miners mine on a chain of length $\geq T$), then all the players will mine the transactions with higher fees first (Lemmas 1 and 2, Appendix 0.C.1). Next, we analyze how the miners behave when the self-penalty mechanism is not employed. Specifically, in Section 0.C.2, we show that without the self-penalty mechanism, each miner that already mined the block $txs_{1}$, will continue to mine this block even when some other player starts to fork this block (Theorem 4). Equipped with this knowledge, the rest of the miners will be discouraged to fork the block $txs_{1}$. Leveraging this result, we then prove in Section 0.C.3, that the majority of the miners will not fork the block $txs_{1}$ in the game without the self-penalty mechanism (Theorem 4). Formally,
{restatable*}
[]thmcontinuemining Let $\Gamma_{F}(\mathcal{S},r+1)$ be a subgame in a state that contains a single chain of length $r\leq T$. The $txs_{p_{1}}$ was not created and a player $N_{j}$ mined $txs_{1}$ in the last round. The player $N_{j}$ will $\mathsf{continue}$ to mine this chain unless $f_{2}-f\geq f_{1}+B$, even when other miners decide to fork the chain with $txs_{1}$ and $\mathsf{continue}$ mining the new subchain created during the fork.
{restatable*}
[]thmnoforks Let $\Gamma_{F}(\mathcal{S},r+1)$ be a subgame in a state $\mathcal{S}$ that contains only a single chain of length $r$ consisting of $r-1$ unrelated transaction sets $txs_{u}$ and one (just mined) $txs_{1}$ (mutually exclusive with $txs_{2}$ with $\mathsf{reward}(txs_{2})>\mathsf{reward}(txs_{1})>\mathsf{reward}(txs_{u})$) mined by some miner $N_{j}$. The $txs_{p_{1}}$ was not created and $r\leq T$. Other miners will not fork $txs_{1}$, unless $f_{2}-f\geq\lambda_{j}(f+B)$, where $\lambda_{j}$ is the mining power of the miner $N_{j}$. These two results let us prove in Section 0.C.4 that in the game with forks where the self-penalty mechanism is not used, the miners will choose to mine $txs_{1}$ in the first round if there exists a player (say player $N_{j}$) with significantly big mining power $>0.01$ that is not encouraged to wait for $txs_{2}$, i.e., $f_{2}-f<\frac{f_{1}-f}{\lambda_{j}}$. The result, formally stated below, is analogous to the result from the game without forks defined in [10]. {restatable*}[]thmtxsone Let $\Gamma_{F}(\mathcal{S},1)$ be a subgame where none of the miners decides to create $txs_{p_{1}}$ before the first round, and the bribing fee is not too high, i.e. $f_{2}-f<10^{-2}(f+B)$. What is more $f_{1}>f$, and if we define $Y=\sum_{j=i:\lambda_{j}>0.01,f_{2}-f<\frac{f_{1}-f}{\lambda_{j}}}^{|N|}\lambda_{j}$, then $T,Y$ are big enough, such that $(1-1.01(1-Y)^{T})>0$. Every miner with $\lambda_{j}>0.01$ will decide to mine $f_{1}$ in the first round. After determining the behaviour of miners when enabling forks (without the self-penalty), we proceed to study the effect of the self-penalty mechanism. We first define a penalty that (i) forces the selected miner $N_{s}$ with mining power $\lambda_{s}$ to fork $txs_{1}$ once it appears on-chain, and (ii) poses a threat strong enough that all the miners avoid mining $txs_{1}$ before the chain length $T$ is reached (Theorem 4).
{restatable*}
[]thmtxstwo Let $\Gamma_{F}(\mathcal{S},2)$ be a subgame where $txs_{p_{1}}$ defined by a player $N_{s}$ with mining power $\lambda_{s}$ was mined in the first round with $P>\lambda_{s}(f+B)$. What is more $f_{2}-f>\frac{f_{1}-f}{\lambda_{s}}$, and $\frac{f+B}{f_{1}+B}>1-\lambda_{s}^{2}$. None of the miners will decide to mine $txs_{1}$ in rounds $2,\ldots,T$. We conclude our analysis with the following result, which states that, overall, using the self-penalty mechanism and mining $txs_{2}$ is more profitable for all players than mining $txs_{1}$ in the first round. We prove Theorem 4 in Section 0.C.6.
{restatable*}
[]thmmain Let $\Gamma_{F}$ be a game with forks that starts with an empty state $\mathcal{S}$, where $10^{-2}(f+B)>f_{2}-f>\frac{2\overline{f}+2(f_{1}-f)}{\lambda_{s}}+\overline{f}$, $\frac{f+B}{f_{1}+B}>1-\lambda_{s}^{2}$, $f_{1}>f$, $f_{p_{2}}>f$, $\lambda_{min}>0.05^{T/2}$ and given $Y=\sum_{j=i:\lambda_{j}>0.01,f_{2}-f<\frac{f_{1}-f}{\lambda_{j}}}^{|N|}\lambda_{j}$, it holds that $(1-1.01(1-(1-Y)^{T/2})>0$. The dominating strategy for all players in the game $\Gamma_{F}$ is is to mine $txs_{p_{1}}$ with $P>\lambda_{s}(f+B)$ created before the first round by the strongest player $N_{s}$ with the mining power $\lambda_{s}$, then mine $txs_{u}$ until round $T$, then $txs_{2}$, $txs_{p_{2}}$, and $txs_{u}$ until the end.
5 Example Evaluation
Using the real-world data analysis of Bitcoin fees and hashpower distribution in major PoW blockchains (see Appendix 0.D), we visualize the improvement our bound $f_{2}-f>\frac{2\overline{f}+2(f_{1}-f)}{\lambda_{s}}+\overline{f}$ brings compared to the previous result from [10], namely $f_{2}-f\geq\frac{f_{1}-f}{\lambda_{min}}$. Additionally, the Theorem 4 requires that $f_{2}-f<10^{-2}(f+B)$ and there exists a player $N_{j}$ with mining power $\lambda_{j}>0.01$ for which $f_{2}-f<\frac{f_{1}-f}{\lambda_{j}}$.
For example, let us assume that $f_{1}-f\approx\overline{f}$, and set $T>110$. Now, since $\lambda_{min}$ can be fairly estimated to be $\lambda_{min}<10^{-12}$, we can see that the attack without forking threats could cost in practice around $10^{12}\overline{f}$. On the other hand, the new bound requires only $f_{2}-f>\frac{2\overline{f}+2(f_{1}-f)}{\lambda_{s}}+\overline{f}$, for $\lambda_{s}\approx 0.2$, this costs around $f_{2}-f>21\overline{f}$. The only condition left is that for some miner with $\lambda_{j}>0.01$, the following condition must hold $f_{2}-f<\frac{f_{1}-f}{\lambda_{j}}$, but the data shows that miners that control approximately $1.5\%-2\%$ of the total mining power usually exist, thus for a miner with mining power $1\%<\lambda_{j}<2\%$ it holds that $f_{2}-f<50\cdot\overline{f}$. In summary, if we take any $f_{2}$ that is larger than $f$ by $21$ up to $50$ times, then the default strategy for all miners is to wait for the bribing transaction.
6 Related Work and Countermeasures
In the landscape of constructing financially stable systems on blockchain [25, 12], our work falls into the class of incentive manipulation attacks. Specifically, to the best of our knowledge, we are the first to combine feather forking attacks [13] with timelock bribing attacks on payment channel networks and to achieve a bribing cost that is approximately only constant times larger than the cost of an average transaction fee.
Incentive manipulation attacks on timelocked puzzles were introduced with the so-called timelock bribing attack [26]. Later, Avarikioti et al. [10] applied timelock bribing attacks in payment channel networks, such as the Lightning Network and Duplex Micropayment Channels, and proposed countermeasures. Our work extends [10], modifying the timelock bribing attack for payment channels to facilitate a miner bribing strategy that incorporates feather forking. As a result, our work reduces the cost of bribing attacks significantly in comparison to [10], i.e., the cost is now inversely proportional to the mining power of the largest miner instead of the smallest miner which results in at least 1000 times smaller bribes.
The bribing strategies for the payment channels are similar in their nature to the bribing strategies for the HTLC mechanism. Perhaps the closest to our work is [34], where the authors introduced a way to bribe HTLCs, leveraging the power of smart contracts and feather forking. The cost of the attack in [34] is, however, proportional to the sum of the fees ($\gtrsim\sum_{i=1}^{T}f\cdot\lambda_{max}$) of all blocks before the deadline $T$. In contrast, we achieve a cost proportional to the cost of fees of a single block ($\gtrsim\frac{f_{1}-f}{\lambda_{S}}$).
Furthermore, MAD-HTLC [32] underlined the vulnerability of HTLCs to bribing attacks, achieving the same attack cost as [10], specifically $\approx\frac{f_{1}-f}{\lambda_{min}}$. MAD-HTLC presupposes that the minimum fraction of mining power controlled by a single user, $\lambda_{min}$, is at least $0.01$, to achieve low bribing costs. This is however an impractical assumption, as the data in Appendix 0.D.2 show that $\lambda_{min}$ can be reasonably estimated to be less than $10^{-12}$, making the bribing attack exceedingly expensive. The reduction of the bribing costs Bribe & Fork achieves in comparison to MAD-HTLC is similar to that of [10] analyzed above.
MAD-HTLC additionally proposed a countermeasure for bribing attacks where miners are allowed to claim the locked coins in the HLTC in case a party misbehaves, similar to [10]. Later, He-HTLC [33] pointed out that MAD-HTLC is susceptible to counter-bribing attacks. In particular, one party may (proactively) collude with the miners to cooperatively steal the coins of the counterparty in the MAD-HTLC construction. He-HTLC also proposed a modification on MAD-HTLC to mitigate the counter-bribing attack: now the coins are partially burned in case of fraud instead of being fully awarded to the miners. Recently, Rapidash [14] revisited the counter-bribing attack and proposed yet another improvement on He-HTLC. These works are orthogonal to ours as the proposed attacks apply only to the specific MAD-HTLC construction and not to Lighting Channels that are the focus of this work. Furthermore, our focus is not on designing countermeasures against timelocked bribing attacks. Instead, we demonstrate how employing feather forking can make timelocked bribing attacks very cheap for the attacker, therefore highlighting the need for robust mitigating strategies.
Nonetheless, it is crucial to acknowledge that the previously mentioned countermeasures can be used to defend against Bribe & Fork– inheriting their respective vulnerabilities. For example, one can employ the mitigation technique for timelocked bribing on the Lighting Network proposed by Avarikioti et al. [10]. In our model, this countermeasure ensures that announcing $txs_{2}$ also involves revealing a secret that anyone can use to claim the money before time $T$. This implies that if $txs_{2}$ is announced in the mining pool before time $T$, all the money to be collected only after time $T$ can be immediately claimed by another party. We assert, without proof, that the same countermeasure mechanism remains effective even in a model that considers forks. Intuitively, the “strong” miner in our analysis does not benefit from waiting for the bribing transaction if it is not announced, thus preventing the creation of the self-penalty transaction. Conversely, if the transaction is announced and the secret is revealed, any (winning) miner could claim the reward money.
Appendix 0.A Details of Bribe & Fork
The Figure 2 contains a diagram depicting the details of our attack. At first, a Lightning Channel is opened with a single Funding transaction, and allows parties $\mathcal{P}_{1},\mathcal{P}_{2}$ to make unlimited number of off-chain state transitions of their funds. Once one of the parties ($\mathcal{P}_{2}$ in our example) decides to publish an old state ($comm_{i}$ in our example), the Game with forks $\Gamma_{F}$ begins. Before round $1$ of the Game, the transactions $tx_{1}$ and $tx_{2}$ are published. Transaction $tx_{1}$ allows $\mathcal{P}_{1}$ to revoke dishonestly committed state. Transaction $tx_{2}$ allows $\mathcal{P}_{2}$ to collect and manage dishonest funds after $T$ rounds. Along with $tx_{1}$ and $tx_{2}$, the bribing transaction $tx_{b}$ and self-penalty transactions $tx_{p_{1}},tx_{p_{2}}$ are published. In round $1$the miners may decide to mine the transaction $tx_{p_{1}}$ that would lock some amount of coins of one of the miners (say $N_{s}$). If so, all miners are threatened that they will be forked once $tx_{1}$ appears on the blockchain until round $T+1$. In round $T+1$, the transaction $tx_{2}$ along with bribing transaction $tx_{b}$ may be published and the selected miner $N_{s}$ may collect back the deposit with the transaction $tx_{p_{2}}$.
Figure 2: The Bribe & Fork attack. The transactions appear on a mempool before the round that they are first shown in the diagram. The green boxes indicate the transactions that should be put on-chain to run a successful Bribe & Fork attack. The grey boxes indicate the transactions that should be published on the mempool before given round. For instance, Spend transaction $tx_{2}$ has to be published on the mempool before round $1$, even though it can not be published on the blockchain until round $T+1$. The arrows going into the boxes indicate the spending conditions of the transactions and the arrows going out of the boxes indicate how the funds of the boxes can be spent.
Appendix 0.B Notation Summary
The Table 1 contains summary of the notation used in the proofs of the theorems.

### v2

@@ 2.5 Bribe & Fork Overview
[…]
Figure 2 contains a diagram depicting the details of our attack. At first, a Lightning Channel is opened with a single funding transaction and allows parties $\mathcal{P}_{1},\mathcal{P}_{2}$ to make an arbitrary number of off-chain state transitions of their funds. Once one of the parties ($\mathcal{P}_{2}$ in our example) decides to publish an old state ($comm_{i}$ in our example) at a chain of length $T_{0}$, the opportunity to manipulate the behaviour of miners’ is opened. Before the chain reaches length $T_{0}+1$, the transactions $tx_{1}$ and $tx_{2}$ are published. Transaction $tx_{1}$ allows $\mathcal{P}_{1}$ to revoke a dishonestly committed state. Transaction $tx_{2}$ allows $\mathcal{P}_{2}$ to collect and manage dishonest funds after $T$ rounds. Along with $tx_{1}$ and $tx_{2}$, the bribing transaction $tx_{b}$ and self-penalty transactions $tx_{p_{1}},tx_{p_{2}}$ are published. On the chain of length $T_{0}$, the miners may decide to mine the transaction $tx_{p_{1}}$ that would lock some amount of coins of one of the miners (say $N_{s}$). If so, all miners are threatened that they will be forked once $tx_{1}$ appears on the blockchain until the chain reaches length $T_{0}+T$. On the chain of length $T_{0}+T$, the transaction $tx_{2}$ along with bribing transaction $tx_{b}$ may be published and the selected miner $N_{s}$ may collect back the deposit with the transaction $tx_{p_{2}}$.
[…]
@@ 3 Our Model
In this section, we introduce our model of PoW mining that takes into account the forks. First, we introduce a set of new notions and then define the game modeling the process of mining.
3.1 Definitions
Let us begin by recalling the conditionally-timelocked output definition from [zeta].
Definition 1 (Conditionally timelocked output - quoted from [zeta])
@@ 3.1 Definitions
For instance, let $\mathcal{S}=\{[(\mathsf{block}_{1},W_{1}),(\mathsf{block}_{2},W_{2})]\}$ be a current state with a single chain $S_{1}$. Then, after a successful fork of $S_{1}$ with $(B_{3},W_{3})$, one gets:
[…]
$[(\mathsf{block}_{1},W_{1}),(\mathsf{block}_{2},W_{2})],$
$[(\mathsf{block}_{1},W_{1}),(\mathsf{block}_{3},W_{3})],$
[…]
Types of Transaction Sets. Each block mined in the game (denoted as $\mathsf{block}$) includes only one of the transaction sets listed below. What is more, the block can be appended to the blockchain, whenever all of the conditions listed below are met:
@@
  • A (bribing) transaction set $txs_{2}$ that contains the transaction $tx_{2}$ that spends money of $txo$ under $cond_{2}$ and a bribing transaction. As long as $txo$ is not spent, this transaction set can be mined on a chain of length $\geq T$. The rest of the transactions in this transaction set are unrelated transactions $tx_{u}$.
@@
Rewards. We assume that a miner, after successfully mining a transaction set $txs_{i}$ on the main chain, gets a reward $\mathsf{reward}(txs_{i})$ equal to $B+f_{i}+P$, where $B$ is a constant block reward and $f_{i}$ is a sum of user fees input by users posting transactions in the transaction set $txs_{i}$. Whenever $txs_{i}$ contains a transaction which locks C coins from miner’s account, we set P to be equal to $-C$. Analogously, when a miner collects $C$ as one of the transactions from $txs_{i}$, we set the parameter $P$ to $C$. The reward for mining a block depends on the number of transactions within the block and their fees. The fee for a more complex transaction is typically higher as it occupies more space in a block. In this respect, we make the following assumptions that correspond to the current Bitcoin values (see Section 5.1 for more details):
@@
Mining Power Distribution. We make the following assumptions (see Section 5.2) about the allocation of the mining power among the players (the $\lambda=(\lambda_{1},\ldots,\lambda_{n})$ tuple):
@@
The Abandon Rule. Let us define the abandon rule $\mathsf{abandon}:\mathbb{S}\rightarrow\mathbb{S}$ that makes a chain the only chain in the game whenever it outruns the length of other chains in the game. That is, for any $S_{i}\in\mathcal{S}$:
$\exists_{S_{j}\in\mathcal{S}}\ len(S_{j})\geq len(S_{i})+1$
then the abandon rule removes $S_{i}$ from the state $\mathcal{S}$.
@@ 3.2 The Game
Finally, we give a description of a game that models the process of Proof-of-Work blockchain mining that takes into account the option to conduct a forking of a block. The game proceeds in rounds, in each round the miners are able to choose whether they want to continue mining on of the chains or they want to fork one of the chains.
@@
When the final round $R$ of the game is over, it finishes in some state $\mathcal{S}$ and rewards are given to the players. By $\mathcal{S}^{*}$ let us denote the longest chain in the state $\mathcal{S}$. Whenever the state has multiple longest chains, $\mathcal{S}^{*}$ denotes the oldest of the longest chains of $\mathcal{S}$. After the final round $R$, in the state $\mathcal{S}$, the reward given to a player $N_{i}$ is:
$\mathsf{reward}_{i}(\mathcal{S})=\sum_{(\mathsf{block},W)\in\mathcal{S}^{*}:W=i}\mathsf{reward}(\mathsf{block}),$
@@
3.3 No Shallow Forks Conjecture
We complement our model with a conjecture that players will not start shallow forks in the game. Shallow forks occur whenever miners fork a transaction that does not explicitly block an opportunity to mine some other transaction in the future.
Conjecture 1 (No shallow forks)
3.4 Notation Summary